Calculate the pH of solution formed on mixing 0.2 M NH4Cl & 0.1 M NH3 .The pOH of ammonia solution is 4.75.?

Grade:11

1 Answers

Arun
6 years ago
Dear Sandip

pOH = pKb + log ([salt]/[base])pKb = 4.75[salt] = 0.2M[base] = 0.1MSubstitute into equation:pOH = 4.75 + log (0.2/0.1)pOH = 4.75 + log 2pOH = 4.75 + 0.30pOH = 5.05pH = 14.00-pOHpH = 14.00-5.05pH = 8.95
 
Regards
Arun (askIITians forum expert)